OpenStudy (anonymous):

solve the equation for x. Round your answer to the nearest hundredth? 8=ln(x+3)

OpenStudy (amistre64):

thats not statistics ....

OpenStudy (anonymous):

I know its something different

OpenStudy (amistre64):

e both sides and subtract the 3

OpenStudy (anonymous):

logs

OpenStudy (anonymous):

can u help me set it up

OpenStudy (amistre64):

8 = ln(e^8 - 3 + 3) 8 = ln(e^8 ) 8 = 8

OpenStudy (anonymous):

z=a^logaz

OpenStudy (amistre64):

\[8=ln(x+3)\] \[e^8=e^{ln(x+3)}\] \[e^8=x+3\] \[e^8=3=x\]

OpenStudy (amistre64):

spose to be a -3 not an =3 lol

OpenStudy (anonymous):

huh

OpenStudy (amistre64):

my typo on the last line ... i hit the "=" key instead of the "-" key .. they just put them too close together